OpenStudy (anonymous):

Convert r^2=a^2sin2theta to cartesian coordinates.

OpenStudy (anonymous):

\[r ^{2}=a ^{2}\sin (2\theta)\]

OpenStudy (psymon):

Start off with using an identity on sin2theta. Double angle formula allows you to rewrite it as 2sin(theta)cos(theta). Now use these conversions to try and help you out. \[r^{2} = x^{2} + y^{2}\] \[x = rcos \theta\] \[y = rsin \theta\]

hartnn (hartnn):

it would be easier to do this before using the formulas r^2. r^2 = 2 a^2 (r sin thetha ) (r cos theta)